OpenStudy (anonymous):

My question is: What is the least common denominator for 1/5, 1/4 and 1/9?

OpenStudy (across):

A good idea would be to factor each number into their prime powers. \(5=2^0\cdot3^0\cdot5^1\) \(4=2^2\cdot3^0\cdot5^0\) \(9=2^0\cdot3^2\cdot5^0\) Since you are trying to find \(\gcd(5,4,9)\), you multiply the minimum of each prime power. \(2^0\cdot3^0\cdot5^0=1\) In other words, \(5,4,9\) are pairwise coprime.

OpenStudy (anonymous):

I see what you are explaining.. But the answer is 180 as the LCD... How can I get to this number the fastest?

OpenStudy (across):

Oh, sorry. I got carried away: what you're trying to find is the \(\text{least common multiple}\). In that case, you multiply the greatest prime powers: \(2^2\cdot3^2\cdot5^1=4\cdot9\cdot5=180\)

OpenStudy (anonymous):

oh yes thats right. I am reviewing my math for the ACT placement test so I am having to refresh my math skills... :( I am not as good as most in math.. Really appreciate your help!